) The supply of paper is described by the following equation:

      Qs = 5P + 10

     The demand is described by:

      Qd = 400 – 10P

      If there is a $10 positive externality per unit of output, what is the socially optimal output?

173 units

183 units

193 units

203 units

The supply of paper is described by the following equation:

      Qs = 5P + 10

     The demand is described by:

      Qd = 400 – 10P

      If there is a $10 negative externality per unit of output, what is the socially optimal output?

67 units

87 units

107 units

127 units

The supply of paper is described by the following equation:

      Qs = 5P + 10

     The demand is described by:

      Qd = 400 – 10P

      The free market price is?

$26

$46

$66

$86

The supply of paper is described by the following equation:

      Qs = 5P + 10

     The demand is described by:

      Qd = 400 – 10P

      The free market output is?

120 units

140 units

160 units

180 units
Answer #1

A) Equilibrium occurs where MPB=MSC

MSC=MPC+ Marginal external benefit= 5(P+10)+10

At Equilibrium,

5(P+10)+10= 400-10P

5P+60=400-10P

340/15= P

22.3=P

Socially optimal quantity= 400-10*23= 173 units.

B.At equilibrium, 5(p-10)+10=400-10p

5p-40=400-10p

P=440/15= $29.3

Socially optimal quantity= 400-10*29.3 = 107

C. Free market price is where Qs=Qd

5p+10=400-10p

Free market price= 390/15=$ 26

D. Free market quantity= 400-10*26= 140 units.